# Open-Vocabulary Calibration for Vision-Language Models

### v2

@@ Abstract
Vision-language models (VLMs) have emerged as formidable tools, showing their strong capability in handling various open-vocabulary tasks in image recognition, text-driven visual content generation, and visual chatbots, to name a few. In recent years, considerable efforts and resources have been devoted to adaptation methods for improving downstream performance of VLMs, particularly on parameter-efficient fine-tuning methods like prompt learning. However, a crucial aspect that has been largely overlooked is the confidence calibration problem in fine-tuned VLMs, which could greatly reduce reliability when deploying such models in the real world. This paper bridges the gap by systematically investigating the confidence calibration problem in the context of prompt learning and reveals that existing calibration methods are insufficient to address the problem, especially in the open-vocabulary setting. To solve the problem, we present a simple and effective approach called Distance-Aware Calibration (DAC), which is based on scaling the temperature using as guidance the distance between predicted text labels and base classes. The experiments with 7 distinct prompt learning methods applied across 11 diverse downstream datasets demonstrate the effectiveness of DAC, which achieves high efficacy without sacrificing the inference speed.
@@ 4.2 Motivation
Lower proximity correlates with higher confidence and ECE.
@@ 4.3 Distance-Aware Calibration
Per analysis above, the predicted class with lower textual proximity can result in higher confidence and worse calibration performance. In this work, we propose a post-hoc method to rectify the predicted confidence level. More specifically, we introduce a textual deviation-based score to scale the temperature value, thereby updating the confidence level to align with the true correctness likelihood.
@@ 5.1 Experimental Setup
Implementation details. For the main results, We use CLIP ( ViT-B/16) (Radford et al., 2021) as the pre-trained VLM throughout our experiments and report results averaged over 3 runs. We fine-tune the model with 16 samples per class in a few-shot setting (Zhou et al., 2022a). For the compared tuning methods, we adopt them from the corresponding official implementation. For our proposed DAC, for simplicity, we set the number of nearest textual feature $K$=5 throughout the paper to calculate TD scores. More details on datasets, pre-trained models, hyperparameters, and implementation details can be found in Appendix A.
@@ 5.2 Results
6 Discussion
@@ 7 Related work
Prompt tuning. Thanks to the rich natural language supervision, pre-trained large vision-language models (Jia et al., 2021; Radford et al., 2021; Yao et al., 2022) can effectively comprehend visual concepts and apply them in downstream tasks (e.g. image classification (Radford et al., 2021; Zhou et al., 2022b; Lu et al., 2022; Naeem et al., 2023), image captioning (Hu et al., 2022; Wang et al., 2023b, a), visual question answering (Parelli et al., 2023)) in a zero-shot manner. Despite VLM’s effectiveness in generalizing new visual concepts, the performance of zero-shot VLM still lags behind the fine-tuned performance on specific downstream tasks (Zhang et al., 2023). To further boost the downstream adaptation of pre-trained VLMs, many parameter-efficient tuning methods like prompt tuning (Zhou et al., 2022b, a; Yao et al., 2023; Zhu et al., 2023; Zang et al., 2022) have been proposed for high efficiency. Recently, several works have explored multimodel tuning (Khattak et al., 2023a, b; Shen et al., 2024), which adapts both vision and language branches of CLIP for better modality alignment. Despite the great success of V-L tuning, their effectiveness on safety-related evaluation is less explored. In this work, we first investigate the calibration performance for fine-tuned VLMs. We present that fine-tuned VLMs generally suffer from miscalibration, and our proposed DAC can effectively mitigate the issue by rectifying the logit scale based on textual information.
@@
Confidence calibration. Confidence calibration has been widely studied to ensure that the confidence levels output by models accurately reflect their true performance. To achieve this, A common practice is to employ post-hoc calibration techniques after model training. Generally, post-hoc calibration can be split into two paradigms: scaling-based methods (Guo et al., 2017; Tomani et al., 2022; Yu et al., 2022; Zhang et al., 2020; Joy et al., 2023; Xiong et al., 2023) and bin-based methods (Zadrozny & Elkan, 2001, 2002; Zhang et al., 2020; Roelofs et al., 2022). For scaling-based calibration, a representative method is temperature scaling (Guo et al., 2017), which learns a single scalar for rescaling the logit. Mix-n-Match (Zhang et al., 2020) leverages ensemble and composition techniques to achieve data efficiency and maintain accuracy in confidence estimates. ATS (Joy et al., 2023) modifies the predicted confidence by per-data-point adaptive temperature. For binning-based calibration, the samples are segmented into various bins according to confidence levels, with each bin undergoing calibration individually. We investigate the open-vocabulary calibration of tuned VLMs. We first show that established post-hoc calibration on base classes can not transfer to new classes. To this end, our DAC fixes the logit scale of prediction via textual deviation-informed score in new classes.
@@ 8 Conclusion
Impact Statements
@@ C.3 Accuracy & Confidence of tuned VLMs
We report the average performance on 11 datasets in Table 11 and 12. Although the classification performance on new classes may be comparable or lower, the average predictive confidence of tuned CLIP is still relatively higher than zero-shot CLIP, which results in a larger ECE. Moreover, we take a close look at the confidence level of MaPLe in new classes. Although the accuracy is comparable with zero-shot CLIP, we can find that the confidence level is greatly improved, which indicates the overconfidence issue.

### v4

# Open-Vocabulary Calibration for Fine-tuned CLIP

@@ Abstract
Vision-language models (VLMs) have emerged as formidable tools, showing their strong capability in handling various open-vocabulary tasks in image recognition, text-driven visual content generation, and visual chatbots, to name a few. In recent years, considerable efforts and resources have been devoted to adaptation methods for improving the downstream performance of VLMs, particularly on parameter-efficient fine-tuning methods like prompt learning. However, a crucial aspect that has been largely overlooked is the confidence calibration problem in fine-tuned VLMs, which could greatly reduce reliability when deploying such models in the real world. This paper bridges the gap by systematically investigating the confidence calibration problem in the context of prompt learning and reveals that existing calibration methods are insufficient to address the problem, especially in the open-vocabulary setting. To solve the problem, we present a simple and effective approach called Distance-Aware Calibration (DAC), which is based on scaling the temperature using as guidance the distance between predicted text labels and base classes. The experiments with 7 distinct prompt learning methods applied across 11 diverse downstream datasets demonstrate the effectiveness of DAC, which achieves high efficacy without sacrificing the inference speed. Our code is available at https://github.com/ml-stat-Sustech/CLIP_Calibration.
@@ 4.2 Motivation
(a) Zero-Shot
(b) CoOp
Lower proximity correlates with higher confidence and ECE.
@@ 4.3 Distance-Aware Calibration
Per the analysis above, the predicted class with lower textual proximity can result in higher confidence and worse calibration performance. In this work, we propose a post-hoc method to rectify the predicted confidence level. More specifically, we introduce a textual deviation-based score to scale the temperature value, thereby updating the confidence level to align with the true correctness likelihood.
@@ 5.1 Experimental Setup
Implementation details. For the main results, We use CLIP ( ViT-B/16) (Radford et al., 2021) as the pre-trained VLM throughout our experiments and report results averaged over 3 runs. We fine-tune the model with 16 samples per class in a few-shot setting (Zhou et al., 2022a). For the compared tuning methods, we adopt them from the corresponding official implementation. For our proposed DAC, for simplicity, we set the number of nearest textual features $K$=5 throughout the paper to calculate TD scores. More details on datasets, pre-trained models, hyperparameters, and implementation details can be found in Appendix A.
@@ 5.2 Results
Table 3: Calibration results of ECE (%) across different confidence levels. $\Delta$ shows the improvement achieved by DAC. Bold numbers denote the top-3 most significant improvements.
Method 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1.0
Conf 0.00 19.24 18.86 15.87 20.42 33.28 30.30 37.84 41.60 18.57
CoOp DAC 0.00 4.95 8.17 11.33 6.51 11.42 24.12 17.41 11.37 -0.94
$\Delta$ 0.00 -14.29 -10.69 -4.54 -13.91 -21.86 -6.18 -20.43 -30.23 -19.51
Conf 0.00 0.00 -12.80 3.90 16.73 10.50 38.07 23.93 19.11 12.13
MaPLe +DAC 0.00 -3.62 -15.32 5.72 6.74 3.12 15.45 6.16 9.29 6.55
$\Delta$ 0.00 -3.62 -2.52 1.82 -10.99 -7.38 -22.62 -17.77 -9.82 -5.58
Conf 0.00 -3.82 0.14 -0.10 4.29 6.31 3.48 8.11 1.23 4.86
ProGrad +DAC 0.00 -0.71 0.03 1.30 -1.32 0.40 -0.65 -0.04 0.44 -0.34
$\Delta$ 0.00 3.11 -0.11 1.40 -5.61 -5.91 -4.13 -8.15 -0.79 -5.20
@@
DAC significantly reduces calibration error, especially for high-confidence predictions. To verify where DAC helps the most over the baseline methods, we analyzed the calibration performance across various confidence levels. Specifically, we measure the gap between confidence and accuracy for each bin on the DTD dataset in Table 3. A negative gap indicates underconfident prediction, and vice versa. The difference between the baseline and our method is presented for easy comparison of improvements in each bin. The results show that DAC improves the calibration performance mainly in those confident predictions (the bins 0.6-0.9). In particular, our method can significantly alleviate the overconfidence issue. For example, in the case of CoOp, our method reduces the gap from 41.60 to 11.37, a direct improvement of 30.23, in the bin with confidence 0.9.
@@
6 Discussion
How does DAC perform in full fine-tuning? To evaluate the flexibility of DAC, we provide additional results with a most recent fine-tuning approach named FLYP(Goyal et al., 2023), where both image and text encoders will be updated. Following the official implementation, we fine-tune the model for 10 epochs with a learning rate of 1e-5.
As is shown in Table 5, DAC successfully reduces the calibration error of models finetuned by FLYP on open-vocabulary classes, which is consistent with the results of prompt-tuning methods. Furthermore, we observe that fine-tuning methods like FLYP also tend to overfit the base classes on several datasets, making it more challenging to calibrate on open-vocabulary classes.
Table 5: Calibration of CoOp and FLYP using ECE (%) across 11 datasets. Bold numbers are significantly superior results.
CoOp FLYP
Dataset Conf +DAC Conf +DAC
Caltech101 4.08 3.17 2.80 2.01
OxfordPets 1.83 1.82 0.73 0.51
StanfordCars 12.50 5.16 19.90 15.62
Flowers102 18.34 10.19 20.58 19.09
Food101 3.83 1.78 5.74 4.13
FGVCAircraft 28.44 17.38 36.52 26.84
SUN397 13.70 4.05 21.72 19.09
DTD 26.86 10.48 12.30 6.49
EuroSAT 12.71 8.62 5.32 5.30
UCF101 19.24 8.67 11.98 8.18
ImageNet 10.69 5.67 8.42 7.19
AVG 13.84 7.00 13.27 10.40
@@ 7 Related work
Prompt tuning. Thanks to the rich natural language supervision, pre-trained large vision-language models (Jia et al., 2021; Radford et al., 2021) can effectively comprehend visual concepts and apply them in downstream tasks (e.g. image classification (Radford et al., 2021; Zhou et al., 2022b; Lu et al., 2022; Naeem et al., 2023), image captioning (Hu et al., 2022; Wang et al., 2023b, a), visual question answering (Parelli et al., 2023)) in a zero-shot manner. Despite VLM’s effectiveness in generalizing new visual concepts, the performance of zero-shot VLM still lags behind the fine-tuned performance on specific downstream tasks (Zhang et al., 2023). To further boost the downstream adaptation of pre-trained VLMs, many parameter-efficient tuning methods like prompt tuning (Zhou et al., 2022b, a; Yao et al., 2023; Zhu et al., 2023; Zang et al., 2022) have been proposed for high efficiency. Recently, several works have explored multimodel tuning (Khattak et al., 2023a, b; Shen et al., 2024), which adapts both vision and language branches of CLIP for better modality alignment. Despite the great success of V-L tuning, their effectiveness on safety-related evaluation is less explored. In this work, we first investigate the calibration performance for fine-tuned VLMs. We present that fine-tuned CLIP generally suffers from miscalibration, and our proposed DAC can effectively mitigate the issue by rectifying the logit scale based on textual information.
Confidence calibration. Confidence calibration has been widely studied to ensure that the confidence levels output by models accurately reflect their true performance. To achieve this, A common practice is to employ post-hoc calibration techniques after model training. Generally, post-hoc calibration can be split into two paradigms: scaling-based methods (Guo et al., 2017; Tomani et al., 2022; Yu et al., 2022; Zhang et al., 2020; Joy et al., 2023; Xiong et al., 2023) and bin-based methods (Zadrozny & Elkan, 2001, 2002; Zhang et al., 2020; Roelofs et al., 2022). For scaling-based calibration, a representative method is temperature scaling (Guo et al., 2017), which learns a single scalar for rescaling the logit. Mix-n-Match (Zhang et al., 2020) leverages ensemble and composition techniques to achieve data efficiency and maintain accuracy in confidence estimates. ATS (Joy et al., 2023) modifies the predicted confidence by per-data-point adaptive temperature. For binning-based calibration, the samples are segmented into various bins according to confidence levels, with each bin undergoing calibration individually. Recently, several works have explored calibration in VLMs (Oh et al., 2023; LeVine et al., 2023). In this work, we focus on the open-vocabulary calibration of fine-tuned VLMs. We first show that established post-hoc calibration on base classes can not transfer to new classes. To this end, our DAC fixes the logit scale of prediction via textual deviation-informed score in new classes.
@@ 8 Conclusion
Acknowledgements
This research is supported by the Shenzhen Fundamental Research Program (Grant No. JCYJ20230807091809020). Bob Zhang is supported by the Science and Technology Development Fund (FDCT), Macao, SAR, under Grant 0028/2023/RIA1. Guoqing Wang is supported in part by the National Natural Science Foundation of China under grant U23B2011, 62102069, U20B2063 and 62220106008, the Key R&D Program of Zhejiang under grant 2024SSYS0091. We gratefully acknowledge the support of the Center for Computational Science and Engineering at the Southern University of Science and Technology for our research.
Impact Statement
@@ C.3 Accuracy & Confidence of tuned VLMs
We report the average performance on 11 datasets in Table 13(b) and 14(b). Although the classification performance on new classes may be comparable or lower, the average predictive confidence of tuned CLIP is still relatively higher than zero-shot CLIP, which results in a larger ECE. Moreover, we take a close look at the confidence level of MaPLe in new classes. Although the accuracy is comparable with zero-shot CLIP, we can find that the confidence level is greatly improved, which indicates the overconfidence issue.